# Pragmatic Curiosity: A Hybrid Learning-Optimization Paradigm via Active Inference

### v1

@@ D.2.1 Plume Field Model and Parameters
General Field and Sensor Parameters:
  • Grid Size: 100 × 100 units.
  • Sensor Size ($a$): 1.0 units.
  • Measurement Time ($\Delta t$): 1.0 seconds.
Source Parameters: Eight potential plume sources were defined for the experiments. Their specific physical parameters are detailed as:
Source $\theta$ $R_{s}$ $\gamma$ $V$ $D$
1 [20.0, 20.0] 100.0 50.0 [0.5, 0.5] 10
2 [30.0, 80.0] 100.0 60.0 [-0.3, 0.2] 15
3 [45.0, 55.0] 15.0 50.0 [0.5, 0.5] 10
4 [50.0, 50.0] 18.0 30.0 [-0.3, 0.2] 15
5 [55.0, 45.0] 16.0 40.0 [0.2, -0.4] 12
6 [48.0, 52.0] 17.0 35.0 [0.1, 0.1] 11
7 [52.0, 55.0] 14.0 45.0 [-0.1, -0.1] 13
8 [52.0, 52.0] 18.0 40.0 [0.1, -0.1] 13
D.2.2 Task-Specific Configurations
Source Localization.
  • Goal: Estimate the unknown location $\theta=[x,y]$ of a single active source.
  • Ground Truth: Source 1 was configured as the single active source.
  • Maximum Sensor Threshold ($y_{max}$): 60.0 hits/second.
  • Hypothesis Space: A discrete grid of potential locations spanning [0,100]×[0,100] with a resolution of 5 units, resulting in a 20 × 20 grid of 400 hypotheses.
  • Degree of curiosity: $\beta=0.5$.
Wind Estimation.
  • Goal: Estimate the unknown wind vector $V=[v_{x},v_{y}]$ for a single source with a known location.
  • Ground Truth: Source 2 was used, with its true wind vector set to [-0.3, 0.2].
  • Maximum Sensor Threshold ($y_{max}$): 60.0 hits/second.
  • Hypothesis Space: A discrete grid of potential wind vectors spanning [-1,1]×[-1,1] with a resolution of 0.1 units, resulting in a 20 × 20 grid of 400 hypotheses.
  • Degree of curiosity: $\beta=1.0$.
Active Source Identification.
  • Goal: Identify the subset of active sources from a set of six potential sources with known locations.
  • Ground Truth: Sources 3-8 were used, with Sources 3, 5, 6, and 8 were set as active.
  • Maximum Sensor Threshold ($y_{max}$): 30.0 hits/second.
  • Hypothesis Space: The set of all possible on/off combinations for the six sources. This is a discrete space with $2^{6}=64$ unique hypotheses.
  • Degree of curiosity: $\beta=5.0$.
D.3 Constrained Active Search
D.3.1 Perception in self driving simulation CARLA
@@
Figure 4: Examples of missed object detection by YOLO due to two reasons considered in perception failure case study in Section D.3. Fig (left to right): example of Failure-1 (distance), Failure-2 (poor light) and Failure-1 and Failure-2 both in one scene (distance and poor light), respectively. Bounding boxes for detected objects (misdetections) shown in yellow (red) with detection confidence numbers. Each scene has two cars and a pedestrian.
D.3.2 LLM-based evaluation for CARLA
@@
D.3.3 Task-Specific Configurations
@@
D.4 Composite Bayesian Optimization
D.4.1 Preference Evaluation
@@
D.4.2 Task-Specific Configurations
@@
  • Ground Truth Preference: $g(y)=-(y-y^{\ast})^{2}$, where $y^{\ast}=[1864.7202,11.8199,0.2904]$.
@@
  • Ground Truth Preference: $g(y)=-(y-y^{\ast})^{2}$, where $y^{\ast}=[25.935,57.612,935.5]$.
@@
  • Ground Truth Preference: $g(y)=a^{\intercal}y$, where a=[1,1,2,1]a=[1,-1,2,1]
@@
D.4.3 Extended Baseline Comparison with BOPE
@@
Figure 5: Extended baseline comparison with BOPE for energy resource allocation. Error bars represent $\pm 1$ standard deviation over 5 seeds.
Figure 5 reports their preference scores over random sampling (RS). It is evident that our active-inference method for hierarchical GPs (HGP) consistently discovers higher-preference regions after a brief initial exploration phase, while BOPE variants are highly sensitive to their stage-wise design choices. BOPE-I, being preference-driven in the first phase, initially attains higher preference values but fails to balance exploration and exploitation, leading to poor convergence in the second half; its performance is also sensitive to the precise switching point. BOPE-II explores first and then optimizes, achieving better final performance than BOPE-I, but the strict separation between exploration and exploitation still yields suboptimal outcomes. BOPE-III mixes both aspects but remains more exploitation-centric, performing better than BOPE-I/II yet still below HGP. BOPE-IV and HGP share the idea of refitting both models at each step, but BOPE-IV converges to a lower-preference solution. In contrast, our acquisition strategy jointly leverages information from both the outcome and preference models at every iteration, leading to higher sample efficiency and more reliable discovery of high-preference regions.
Appendix E Use of Large Language Models
We use LLM to revise and polish drafts of the introduction, abstract, conclusion, and limitations sections to enhance clarity, flow, and accessibility for a broader audience.
We also use LLM as an evaluator for the experiments of failure discovery in CARLA, which is detailed in Appendix D.3.

### v2

@@ C.2.1 Plume Field Model
C.2.2 Task 1: Source Response Localization
The first task asks the agent to localize a plume source well enough to dispatch a response team near it. The true source is located at $(35,65)$. The latent hypotheses $\theta$ consist of source-location hypotheses on a grid over the $100\times 100$ domain with spacing $5$, together with source-strength multipliers
$[0.4,0.7,1.0,1.6].$
The response actions form a coarser grid with spacing $20$. The prior is intentionally multimodal, with modes near $(75,25)$ and $(35,65)$, so the agent must resolve a decision-relevant ambiguity rather than simply exploit a unimodal belief.
For a response action $a$ and hypothesis $\theta$, let $\ell_{\theta}$ denote the source location specified by $\theta$. The downstream loss is the normalized squared response distance,
$L_{\mathrm{loc}}(a,\theta)=\frac{\|a-\ell_{\theta}\|_{2}^{2}}{100^{2}+100^{2}}.$
Given posterior $q_{t}(\theta)=p(\theta\mid\mathcal{D}_{t})$, the Bayes risk is
$BR(p_{t})=\min_{a}\sum_{\theta}q_{t}(\theta)L_{\mathrm{loc}}(a,\theta),$
and the Bayes action is
$a_{t}^{\star}=\arg\min_{a}\sum_{\theta}q_{t}(\theta)L_{\mathrm{loc}}(a,\theta).$
For dynamic curiosity schedule, the default dynamic settings are
$q=0.75,\quad k_{\beta}=1,\quad\varepsilon_{I}=10^{-8},\quad\varepsilon_{U}=10^{-8},\quad\beta_{\min}=10^{-3},\quad\beta_{\max}=10,\quad U^{\star}=0.$
We report:
  • Bayes risk: $BR(p_{t})$.
  • Response loss: realized loss of the Bayes action under the true source,
    $L_{\mathrm{loc}}(a_{t}^{\star},\theta_{\mathrm{true}}).$
  • Response success: whether the selected response action lies within one response-grid spacing of the true source,
    $\mathbf{1}\{\|a_{t}^{\star}-\ell_{\theta_{\mathrm{true}}}\|_{2}\leq 20\}.$
  • Parameter error: Euclidean error between the posterior MAP source location and the true source location.
C.2.3 Task 2: Consequence-Weighted Dispatch
The second task uses the same response-decision structure, but errors in high-consequence regions are more costly. The true source is located at $(68,72)$. The prior has modes near $(34,66)$, $(68,72)$, and $(76,24)$, and the high-consequence region is centered near the true source.
Let $C(\ell_{\theta})$ denote the consequence weight associated with the source location under hypothesis $\theta$. In the implementation, this weight is defined by a baseline term plus Gaussian consequence regions,
$C(\ell_{\theta})=C_{0}+\sum_{j}w_{j}\exp\left(-\frac{\|\ell_{\theta}-c_{j}\|_{2}^{2}}{2\sigma_{j}^{2}}\right),$
with an additional source-strength multiplier. The consequence-weighted downstream loss is
$L_{\mathrm{cw}}(a,\theta)=C(\ell_{\theta})\frac{\|a-\ell_{\theta}\|_{2}^{2}}{100^{2}+100^{2}}.$
Given posterior $q_{t}(\theta)$, the Bayes risk is
$BR(p_{t})=\min_{a}\sum_{\theta}q_{t}(\theta)L_{\mathrm{cw}}(a,\theta).$
For dynamic curiosity schedule, the default dynamic settings are
$q=0.75,\quad k_{\beta}=1,\quad\varepsilon_{I}=10^{-8},\quad\varepsilon_{U}=10^{-8},\quad\beta_{\min}=10^{-3},\quad\beta_{\max}=10,\quad U^{\star}=0.$
We report:
  • Bayes risk: posterior expected consequence-weighted loss of the Bayes action.
  • Response loss: realized consequence-weighted response loss under the true source.
  • Response success: the same geometric success criterion as in source response localization.
  • Parameter error: Euclidean MAP source-location error.
C.2.4 Task 3: Active Source Prioritization
The third task models limited repair or inspection resources. There are six possible sources, and the true active sources are indices $[4,5]$. The repair budget is $k=2$. The source consequence weights are
$[70,65,5,5,180,170],$
and the prior activity probabilities are
$[0.82,0.80,0.40,0.40,0.15,0.15].$
This creates a decision tension: the prior favors high-probability decoys, while the true active sources have high consequence weights but low prior activity.
The posterior is over nonempty active-source subsets $\theta$. Let
$z_{i}(\theta)\in\{0,1\}$
indicate whether source $i$ is active under hypothesis $\theta$. The posterior marginal activity probability of source $i$ is
$p_{t,i}=\sum_{\theta}q_{t}(\theta)z_{i}(\theta).$
With zero false-positive cost, the expected missed-source risk before selecting repairs is
$\sum_{i}w_{i}p_{t,i}.$
If the selected repair set is $S$, its posterior expected benefit is
$\sum_{i\in S}w_{i}p_{t,i}.$
The Bayes action selects the top $k$ sources by $w_{i}p_{t,i}$, and the Bayes risk is
$BR(p_{t})=\sum_{i}w_{i}p_{t,i}-\max_{|S|=k}\sum_{i\in S}w_{i}p_{t,i}.$
For dynamic curiosity schedule, the default dynamic settings are
$q=0.9,\quad k_{\beta}=2,\quad\varepsilon_{I}=10^{-8},\quad\varepsilon_{U}=10^{-8},\quad\beta_{\min}=1,\quad\beta_{\max}=10,\quad U^{\star}=0.$
We report:
  • Bayes risk: posterior expected weighted risk left unrepaired after selecting $k$ sources.
  • Missed-source risk: realized total weight of true active sources not selected,
    $\sum_{i\in A_{\mathrm{true}}\setminus S_{t}}w_{i}.$
  • Top-$k$ recall:
    $\frac{|S_{t}\cap A_{\mathrm{true}}|}{\min(k,|A_{\mathrm{true}}|)}.$
  • Weighted top-$k$ recall:
    $\frac{\sum_{i\in S_{t}\cap A_{\mathrm{true}}}w_{i}}{\sum_{i\in A_{\mathrm{true}}}w_{i}}.$
  • Parameter error: symmetric-difference size between the MAP active-source set and the true active-source set.
C.2.5 Baselines
We compare the following methods.
Random.
Random selects a sensor query uniformly or according to the predefined random sampling protocol over the candidate set.
Pure information gain.
EIG selects
$x_{t}=\arg\max_{x\in\mathcal{X}_{t}^{\mathrm{cand}}}\widehat{I}_{t}(x),$
and therefore learns the latent environment without regard to downstream decision consequence.
Decision-greedy.
Greedy selects
$x_{t}=\arg\max_{x\in\mathcal{X}_{t}^{\mathrm{cand}}}\widehat{\Delta BR}_{t}(x),$
and therefore maximizes immediate expected downstream Bayes-risk reduction without explicitly valuing epistemic clarification.
Fixed-$\beta$ PraC.
Fixed PraC uses
$\alpha_{t}(x)=\widehat{\Delta BR}_{t}(x)+5\,\widehat{I}_{t}(x).$
Dynamic-$\beta$ PraC.
Dynamic PraC uses the same acquisition but computes $\beta_{t}$ adaptively from the feedforward–feedback schedule described above.
C.2.6 Analysis of Decision-Oriented Monitoring Results
The decision-oriented plume-monitoring experiments are designed to test whether PraC learns environmental distinctions that matter for downstream action, rather than merely reducing parameter uncertainty. The detailed results are reported in Table 2.
Across the three tasks, the results show three complementary behaviors: dynamic PraC improves downstream Bayes-risk reduction in the response-localization and source-prioritization tasks; decision-greedy can be competitive when posterior risk is already well aligned with the downstream loss; and pure information gain can fail when uncertainty reduction is not sufficiently targeted toward the decision-relevant ambiguity.
Source response localization.
In source response localization, dynamic PraC achieves the lowest final Bayes risk, $0.0025\pm 1.59\times 10^{-12}$, and also obtains zero parameter error. All non-random adaptive methods achieve the best realized response loss and response success, indicating that the response decision itself is relatively easy once the posterior identifies a response-equivalent source region. However, the Bayes-risk and parameter-error results reveal a sharper distinction among methods. Dynamic PraC resolves the latent ambiguity most completely, whereas fixed PraC and EIG retain residual MAP source-location error, and Greedy retains larger parameter error. This suggests that, in this task, downstream decision improvement and epistemic clarification are aligned, but dynamic curiosity still improves the posterior quality beyond what is required for merely selecting a successful response action.
Consequence-weighted dispatch.
The consequence-weighted dispatch task is more diagnostic because the posterior Bayes risk and the realized response loss no longer rank methods identically. Greedy achieves the lowest final Bayes risk, $0.0254\pm 0.00538$, because it directly optimizes the one-step expected reduction in posterior risk. However, dynamic PraC achieves the best realized response loss, $0.0347\pm 0.00771$, and the lowest parameter error, $7.62\pm 5.56$. This gap is important: posterior Bayes risk measures expected loss under the agent’s current belief, whereas realized response loss evaluates the downstream decision under the true latent hypothesis. A purely greedy policy can therefore reduce posterior risk quickly while still failing to collect information that would improve the realized decision under the true source. Dynamic PraC sometimes sacrifices immediate posterior-risk reduction to clarify the environmental hypothesis, which improves the eventual response decision in the realized environment.
Active source prioritization.
Active source prioritization creates a different form of decision ambiguity. The prior assigns high activity probabilities to some lower-consequence decoy sources, while the true active sources have lower prior probability but much larger consequence weights. Dynamic PraC achieves the lowest final Bayes risk, $2.86\pm 2.72$, improving over fixed PraC, Greedy, EIG, and Random. Dynamic and fixed PraC tie on missed-source risk, top-$k$ recall, and weighted top-$k$ recall, showing that both PraC variants reliably identify the high-consequence sources needed for the repair decision. By contrast, EIG and Random perform substantially worse on missed-source risk and weighted recall. This supports the central point of the decision-oriented formulation: information is useful only when it helps resolve the ambiguity that affects the downstream repair set. Pure information gain may reduce uncertainty over the active-source subset, but it does not necessarily prioritize distinctions that change the top-$k$ consequence-weighted decision.
Comparison with pure information gain.
EIG is a strong baseline when learning the latent hypothesis is well aligned with the downstream decision. This explains why it performs competitively in source response localization and matches fixed PraC on several metrics. However, EIG lacks an explicit mechanism for distinguishing decision-relevant uncertainty from decision-irrelevant uncertainty. In consequence-weighted dispatch and active source prioritization, this limitation becomes more visible: EIG can collect information about the environment without adequately targeting the distinctions that reduce consequence-weighted response loss or missed-source risk. PraC addresses this by coupling information gain with expected Bayes-risk reduction.
Comparison with decision-greedy.
The decision-greedy baseline isolates the pragmatic term by choosing queries that maximize $\widehat{\Delta BR}_{t}(x)$ without any explicit epistemic value. Its strong Bayes-risk performance in consequence-weighted dispatch shows that myopic posterior-risk reduction can be effective when the current belief already points toward useful response decisions. However, Greedy is more vulnerable when the current posterior is brittle or biased by the prior. In active source prioritization, Greedy is substantially worse than dynamic PraC in final Bayes risk, even though it remains competitive on some top-$k$ metrics. This indicates that myopic decision improvement may not be sufficient when the downstream action depends on resolving low-prior but high-consequence hypotheses.
Why Bayes risk and realized loss can disagree.
The distinction between Bayes risk and realized downstream loss is essential for interpreting these experiments. Bayes risk is the expected loss under the agent’s posterior belief, while realized loss is evaluated under the true latent hypothesis. A method can reduce Bayes risk by becoming confident under its current posterior, even if the posterior remains biased in a way that hurts realized performance. This is why Greedy can obtain the lowest final Bayes risk in consequence-weighted dispatch while dynamic PraC obtains the best realized response loss and parameter error. PraC’s epistemic term helps correct the posterior before committing too strongly to the current decision surrogate.
Overall, the three tasks show that PraC is most beneficial when downstream action depends on resolving specific decision-relevant ambiguity. When the pragmatic objective and epistemic objective are aligned, PraC remains competitive with EIG and Greedy. When they diverge, dynamic PraC provides a mechanism for deciding when information is worth acquiring because it improves downstream action. This supports the role of PraC as a hybrid learning-and-optimization framework: the agent does not learn the environment for its own sake, nor does it merely optimize the current posterior decision; it learns the environmental distinctions that make better downstream action possible.
Table 2: Full final performance metrics for decision-oriented plume monitoring over 20 seeds. Lower is better for Bayes risk, response loss, parameter error, and missed-source risk. Higher is better for response success, top-$k$ recall, and weighted top-$k$ recall. Bold indicates the best or tied-best mean within each task and metric. Dyn. PraC denotes dynamic-$\beta_{t}$ PraC; Fixed PraC denotes fixed-$\beta$ PraC with $\beta=5$.
Source Response Localization
Metric Dyn. PraC Fixed PraC Greedy EIG Random
Bayes risk $\downarrow$ $\mathbf{0.0025\pm 1.59{\times}10^{-12}}$ $00377\pm 0.00254 0.0$$0463\pm 0.000765 0.0$$0377\pm 0.00254 0.0$$0124\pm 0.00495$
Response loss $\downarrow$ $\mathbf{0.0025\pm 0}$ $\mathbf{0.0025\pm 0}$ $\mathbf{0.0025\pm 0}$ $\mathbf{0.0025\pm 0}$ $0.0125\pm 0.00632$
Response success $\uparrow$ $\mathbf{1.00\pm 0}$ $\mathbf{1.00\pm 0}$ $\mathbf{1.00\pm 0}$ $\mathbf{1.00\pm 0}$ $0.80\pm 0.40$
Parameter error $\downarrow$ $\mathbf{0.00\pm 0}$ $2.00\pm 4.00$ $5.24\pm 5.25$ $2.00\pm 4.00$ $8.61\pm 5.09$
Consequence-Weighted Dispatch
Metric Dyn. PraC Fixed PraC Greedy EIG Random
[…]
Response loss $\downarrow$ $\mathbf{0.0347\pm 0.00771}$ $0424\pm 0.00945 0.0$$424\pm 0.00945 0.0$$424\pm 0.00945 0.0$$0463\pm 0.0144$
Response success $\uparrow$ $\mathbf{1.00\pm 0}$ $\mathbf{1.00\pm 0}$ $\mathbf{1.00\pm 0}$ $\mathbf{1.00\pm 0}$ $\mathbf{1.00\pm 0}$
Parameter error $\downarrow$ $\mathbf{7.62\pm 5.56}$ $8.86\pm 6.60$ $04\pm 3.06 8.8$$86\pm 6.60$ $17.8\pm 9.72$
Active Source Prioritization
Metric Dyn. PraC Fixed PraC Greedy EIG Random
Bayes risk $\downarrow$ $\mathbf{2.86\pm 2.72}$ $3.59\pm 2.70$ $14.0\pm 10.6$ $23.3\pm 19.1$ $34.5\pm 17.0$
Missed-source risk $\downarrow$ $\mathbf{34.0\pm 68.0}$ $\mathbf{34.0\pm 68.0}$ $36.0\pm 72.0$ $68.0\pm 83.3$ $316\pm 68.0$
Top-$k$ recall $\uparrow$ $\mathbf{0.90\pm 0.20}$ $\mathbf{0.90\pm 0.20}$ $\mathbf{0.90\pm 0.20}$ $0.80\pm 0.245$ $0.10\pm 0.20$
Weighted top-$k$ recall $\uparrow$ $\mathbf{0.903\pm 0.194}$ $\mathbf{0.903\pm 0.194}$ $0.897\pm 0.206$ $806\pm 0.238 0.0$$0971\pm 0.194$
C.2.7 Diagnostics of Dynamic Curiosity
In addition to the main performance curves, we report scheduler diagnostics to verify that dynamic curiosity behaves as intended. The diagnostic plots include:
  • $\beta_{t}$, the scheduled curiosity coefficient;
  • $\beta_{t}^{\mathrm{ff}}$, the feedforward exchange-rate scale;
  • $\beta_{t}^{\mathrm{fb}}$, the feedback activation factor;
  • $\beta_{t}\widehat{I}_{t}(x_{t})$, the selected effective epistemic pressure.
Figure 5: Diagnostics of dynamic curiosity scheduler. Shaded regions show mean $\pm 1$ standard error over 20 seeds.
These diagnostics distinguish three regimes. When decision-relevant uncertainty is high, the feedback activation keeps curiosity active. When useful information is costly relative to downstream Bayes-risk reduction, the feedforward scale increases. When the decision-symbol posterior concentrates, the feedback activation suppresses curiosity even if raw information-gain ratios become numerically large. This prevents late-stage ratio explosions and makes the dynamic schedule interpretable.
Role of dynamic curiosity.
The dynamic schedule gives $\beta_{t}$ an operational meaning as an exchange rate between epistemic value and downstream Bayes-risk reduction. The feedforward scale estimates how large curiosity must be for information gain to compete with pragmatic improvement, while the feedback activation suppresses curiosity as decision-relevant uncertainty decreases. This prevents $\beta_{t}$ from acting as a fixed exploration knob. Instead, the relevant quantity is the effective epistemic pressure $\beta_{t}\widehat{I}_{t}(x)$. The empirical results are consistent with this interpretation: dynamic PraC improves posterior quality in source response localization, improves realized downstream loss in consequence-weighted dispatch, and achieves the lowest Bayes risk in active source prioritization.
C.3 Targeted Active Search
C.3.1 Perception in self driving simulation CARLA
@@
Figure 6: Examples of missed object detection by YOLO due to two reasons considered in perception failure case study in Section C.3. Fig (left to right): example of Failure-1 (distance), Failure-2 (poor light) and Failure-1 and Failure-2 both in one scene (distance and poor light), respectively. Bounding boxes for detected objects (misdetections) shown in yellow (red) with detection confidence numbers. Each scene has two cars and a pedestrian.
C.3.2 LLM-based evaluation for CARLA
@@
C.3.3 Task-Specific Configurations
@@
C.4 Composite Bayesian Optimization
C.4.1 Preference Evaluation
@@
C.4.2 Task-Specific Configurations
@@
  • Ground Truth: $g(y)=(y-y^{\ast})^{2}$, where $y^{\ast}=[1864.7202,11.8199,0.2904]$.
@@
  • Ground Truth: $g(y)=(y-y^{\ast})^{2}$, where $y^{\ast}=[25.935,57.612,935.5]$.
@@
  • Ground Truth Preference: $g(y)=a^{\intercal}y$, where a=[1,1,2,1]a=[-1,1,-2,-1]
@@
C.4.3 Comparison with BOPE
@@
Figure 7: Extended baseline comparison with BOPE for energy allocation. Error bars represent $\pm 1$ standard deviation over 20 seeds.
Figure 7 reports their preference scores over random sampling (RS). It is evident that our method (PraC) consistently discovers higher-preference regions after a brief initial exploration phase, while BOPE variants are highly sensitive to their stage-wise design choices. BOPE-I, being preference-driven in the first phase, initially attains higher preference values but fails to balance exploration and exploitation, leading to poor convergence in the second half; its performance is also sensitive to the precise switching point. BOPE-II explores first and then optimizes, achieving better final performance than BOPE-I, but the strict separation between exploration and exploitation still yields suboptimal outcomes. BOPE-III mixes both aspects but remains more exploitation-centric, performing better than BOPE-I/II yet still below PraC. BOPE-IV and PraC share the idea of refitting both models at each step, but BOPE-IV converges to a lower-preference solution. In contrast, our acquisition strategy jointly leverages information from both the outcome and preference models at every iteration, leading to higher sample efficiency and more reliable discovery of high-preference regions.